# Human-in-the-Loop Feature Selection Using Interpretable Kolmogorov-Arnold Network-based Double Deep Q-Network

### v1

@@ 22             Store $(s,a,s^{\prime},r)$ in $\mathcal{B}$;
23
[…]
@@ 42      Adjust learning rate with scheduler;
43
44Return $Q_{\theta}$ and collected metrics (accuracy, loss, feedback cost history);
@@
III-B Double Deep Q-Network Architecture
@@
III-C KAN
In our approach, the KAN-based architecture is structured with widths of [64, 8, 10] for the input, hidden, and output layers. Both the Q-network and target network use this compact configuration to capture complex feature interactions efficiently. Each KAN layer combines a spline function and a residual basis function, formulated as:
@@ III-C KAN
where $b(x)=silu(x)$ serves as a non-linear basis function, while $spline(x)=\sum_{i}c_{i}B_{i}(x)$ utilizes trainable B-splines for flexible approximation. With spline order $k=3$ over a specified grid, this layered structure dynamically adjusts to feature relevance throughout training. For optimization, $w_{b}$ and $w_{s}$ are initialized carefully: $w_{b}$ uses Xavier initialization, ensuring balanced layer activation scales, while $w_{s}$ is set to 1, with the initial spline function close to zero. The model’s spline grids adapt to changing input activations, extending KAN’s effective region for learning. By leveraging these optimizations and not training biases or spline parameters directly, the KAN achieves significant performance gains with reduced neurons in the hidden layer—showing a parameter-efficient architecture that improves interpretability.
III-D MLP
The MLP-based architecture provides a more traditional setup for the DDQN framework, using a straightforward, fully connected design for both the Q-network and the target network. This configuration includes 2 linear layers: a 64-neuron initial layer to map the input features to a hidden space defined by a 32-neuron network width, followed by a ReLU activation, and a 10-neuron final linear layer to produce Q-values. The MLP requires a larger hidden layer with 32 neurons to approximate the complex mappings found in datasets, making it parameter-heavy relative to the KAN-based DDQN.
III-E Training Procedure: KAN-based and MLP-based DDQN
The training procedure for the DDQN is illustrated in Algorithm 2, which outlines the main steps taken during each epoch. The algorithm initializes the environment and Q-networks and iteratively processes batches of experiences from a replay buffer $\mathcal{B}$. Each experience consists of the current state $s$, action $a$, reward $r$, and the next state $s^{\prime}$.
We employ two training algorithms to optimize our Q-network, each tailored to different model architectures: (1) a KAN-based training procedure that utilizes regularization for interpretability and stability and (2) a standard MLP-based training procedure. Each training procedure integrates feedback cost when feature selection is enabled, promoting a minimalistic and interpretable feature set.
III-E1 Temporal-Difference Target Computation
@@
[…]
@@
III-E2 KAN-based Training with Regularization
@@
[…]
@@
In addition, an entropy term is introduced to reduce overconfidence in the predictions. The entropy $S(\Phi)$ for the KAN layer is given by:
@@
$\mathcal{L}_{\text{KAN}}=\mathcal{L}_{\text{TD}}+\lambda\mathcal{R}=L_{\delta}\left(Q_{\theta}(s,a),\textit{TD\_target}\right)+\lambda\left(\mathcal{R}_{\text{L1+Entropy}}+\mathcal{R}_{\text{Spline}}\right)$ (28)
III-E3 Feedback Cost Integration for Feature Selection
@@
III-E4 MLP-based Training
For the MLP-based training, we use the same TD target computation and $SmoothL1$ loss but omit the KAN-specific regularization terms. Like KAN-based training, feature selection introduces an additional feedback cost term to this loss: $\mathcal{L}_{\text{MLP, total}}=L_{\delta}\left(Q_{\theta}(s,a),\textit{TD\_target}\right)+\alpha\times\mathcal{L}_{\text{feedback}}$.
III-E5 Optimization and Gradient Clipping
@@
III-F Integration of HITL Feedback in DDQN
@@
III-F1 Feature Selection Adjustment Using HITL Feedback
@@
III-F2 Action-Selection Policy with HITL Feedback
@@
III-F3 Iterative Learning Process
@@
IV Experimental Design
IV-A Dataset Preparation
We used a benchmark environment to standardize data preparation, feedback simulation, and visualizations across two datasets—MNIST and FashionMNIST. Each dataset was resized to a standardized input dimension of $8\times 8$ pixels, with normalization applied based on empirically computed means and standard deviations specific to each dataset to aid in training convergence. For MNIST, we used a mean ($\mu$) of 0.1307 and a standard deviation ($\sigma_{SD}$) of 0.3081, while for FashionMNIST, the $\mu$ and $\sigma_{SD}$ were both set to 0.5. This normalization process adjusts pixel intensity $\mathbf{I}$ via the formula:
@@ IV-A Dataset Preparation
IV-B Simulated Feedback Mechanism
@@ IV-C Resources Used
IV-D Experimental Configurations
For the experiments conducted in this study, we employed two configurations for the MLP and KAN models, both set to a batch size of 128 and trained over 100 epochs. The configurations utilized a learning rate $1\times 10^{-3}$, a weight decay $1\times 10^{-4}$, and a discount factor ($\gamma$) of 0.99. The MLP configuration featured a width of 32 and an output size of 10 classes, while the KAN configuration used a width of 8 and a grid size of 3. Both models share a common input size of 64 features, a buffer size 100,000, and a target update interval of 10. They incorporated warm-up episodes of 2 and utilized a beta distribution (see Table I) for their stochastic feature selection processes, with an initial $\tau$ value set at 1.0. We strategically designed the architecture of our model by selecting the appropriate combination of convolutional layers and filters to ensure that the resulting selected feature map retains the same shape as the input and feedback size (64 for $8\times 8$ images). This choice is crucial, as it directly allows the simulated feedback (similar to input size) to correspond to the feature map dimensions, facilitating effective integration during the feature selection process. In this case, for 1, 2, and 3 convolutional layers, we must select 2, 4, and 8 filter sizes to maintain the desired selected feature shape. The means and standard deviations for training and test accuracies are computed over a 10-epoch moving window to capture stability and uncertainty trends throughout training.
Figure 2: Comparison of KAN and MLP-based DDQN on MNIST (left) and FashionMNIST (right) involving (a) training accuracy per epoch and (b) confusion matrix.
Figure 3: (a) Pruned KAN architecture for MNIST, (b) feature importance plot obtained from the symbolic formula of the trained KAN agent, (c) training steps of our KAN agent for MNIST dataset.
Figure 4: Activation functions for middle neurons in KAN agent’s hidden layer on MNIST dataset. Each subplot shows how a specific neuron transforms inputs via the spline activation, illustrating various activation behaviors across the layer.
TABLE I: Performance of MLP-DDQN for Different Distributions on MNIST and Fashion MNIST Datasets. Bold indicates the best performance.
@@ IV-D Experimental Configurations
TABLE II: Performance Metrics without Feature Selection on MNIST and FashionMNIST Datasets Using KAN and MLP Models.
@@
maP maR maF1 Train Accuracy Test Accuracy
KAN MNIST 0.57 0.57 0.56 0.56 $\pm$ 0.03 0.58 $\pm$ 0.01 5:10:50
MLP MNIST 0.45 0.51 0.47 0.47 $\pm$ 0.02 0.52 $\pm$ 0.00 10:41
KAN FashionMNIST 0.63 0.65 0.63 0.67 $\pm$ 0.02 0.64 $\pm$ 0.00 5:10:32
MLP FashionMNIST 0.55 0.60 0.57 0.59 $\pm$ 0.02 0.59 $\pm$ 0.00 10:24
TABLE III: Performance Comparison of KAN and MLP Models on MNIST and FashionMNIST with Feature Selection
@@
TABLE IV: Impact of KAN Parameter Adjustments on MNIST and FashionMNIST Performance
@@ V Results
Our results reveal that KAN-DDQN outperforms MLP-DDQN in all cases on both MNIST and FashionMNIST datasets (see Fig. 2a and 2b) with 4 times fewer parameters and only $8\times 8$ sized images.
V-A Effect of Distribution Type in Feature Selection
@@
Table I demonstrates significant variability in MLP-DDQN performance across different distributions for both MNIST and FashionMNIST datasets. We conducted this experiment to choose the best distribution that can mimic the human-like feature selection process. The beta distribution consistently yields the highest train and test accuracies on MNIST (79% and 78%, respectively), prompting us to conduct the remaining experiments with this distribution. Conversely, the Gumbel Softmax, Dirichlet, multinomial, Cauchy, and Laplace distributions perform poorly, with test accuracies as low as 12%. Notably, the uniform distribution also shows competitive results, particularly on the MNIST dataset, suggesting it can serve as a viable alternative for feature selection.
V-B Effect of Feature Selection in Performance
V-B1 Performance With Feature Selection
@@
Table III illustrates the performance of the KAN-DDQN and standard MLP-DDQN using feature selection across MNIST and FashionMNIST datasets. In all cases, the hidden layer width was set to 32 neurons for the MLP and 8 for the KAN, reflecting the distinct architectures of each model. The KAN model consistently outperformed the MLP across all configurations, particularly on the MNIST dataset, achieving a test accuracy of 93% with a single convolutional layer, two filters, a width of 8, and a grid size of 3. For FashionMNIST, KAN similarly demonstrated superior test performance, reaching 83% accuracy under the same configuration, compared to MLP’s 74% accuracy. Across both datasets, KAN exhibited higher mean average precision (maP), recall (maR), and F1-score (maF1). For example, in the single-layer configuration on MNIST, KAN achieved scores of 0.92 for each metric, whereas MLP scored 0.84. The training time for the KAN model was generally longer because of its more intricate architecture, with the best-performing KAN model on MNIST requiring 4 hours and 35 minutes, compared with only 8 minutes for MLP (see Table III). Despite the increased computation, the accuracy gains suggest that KAN is efficient for applications where performance outweighs training time constraints. Notably, with the addition of a convolutional layer and increased filter size (e.g., 2 layers, 4 filters), KAN’s performance slightly declined, particularly on FashionMNIST, where test accuracy was 70%. This indicates a potential trade-off between model complexity and generalization, suggesting that simpler KAN configurations, particularly with a lower hidden layer width, may better retain discriminative power for smaller datasets.
V-B2 Performance Without Feature Selection
@@
Table II provides a comparative view of model performance without feature selection, highlighting the impact of excluding feature selection on both models’ performance. Without feature selection, both KAN and MLP demonstrate considerably poor performance. Here, the KAN model achieves a notable decrease in test accuracy and maF1 compared with the MLP model. Specifically, on MNIST, KAN reaches a test accuracy of 58% with maF1 of 0.56, surpassing the MLP model’s 52% test accuracy and 0.47 maF1. On FashionMNIST, the KAN model continues to outperform, achieving a 64% test accuracy and 0.63 maF1, compared with MLP’s 59% accuracy and 0.57 maF1.
V-C Effect of KAN Parameters in Performance
@@
Table IV shows how KAN performance varies with parameter adjustments, focusing on width (number of neurons in hidden layer) and grid size. Doubling the width of the feature maps from 8 to 16 yielded improvements in accuracy, especially on MNIST, where the test accuracy rose from 72% to 75%. However, this increase came at the expense of significantly longer training times, suggesting a trade-off between accuracy and computational efficiency. On FashionMNIST, similar parameter adjustments yielded more moderate gains. For example, increasing the grid size to 3 and width to 16 raised test accuracy slightly from 71% to 72%. Moreover, altering the grid size from 2 to 3 produces varied outcomes: it preserves similar test accuracy in MNIST (0.72 to 0.73) but slightly decreases performance in FashionMNIST (0.71 to 0.70). This variation may stem from applying the B-spline in higher dimensions, warranting further exploration. This pattern suggests that KAN’s parameter sensitivity is dataset-dependent, with MNIST benefiting more from only width increases. Training time increased substantially with larger configurations. On MNIST, the KAN model with a grid size of 3 and width of 16 took over 10 hours, while the smaller grid size of 2 and width of 8 configurations was completed in 6 hours. Thus, reducing width and grid size may be advantageous for time-sensitive applications, especially where marginal accuracy improvements are not critical. So, larger hidden layer widths improve KAN’s accuracy but at the cost of increased training time. Balancing grid size and width for applications prioritizing efficiency may offer an optimal trade-off without sacrificing significant accuracy.
V-D Interpretability of KAN-DDQN

### v2

@@ 24     Store $(s,a,s^{\prime},r)$ in $\mathcal{B}$;
25
26    if buffer $\mathcal{B}$ is ready (size $\textit{batch\_size}$) then
@@ 49 Adjust learning rate with scheduler;
50
51 end for
52Return $Q_{\theta}$ and collected metrics (accuracy, loss, feedback cost history);
@@
3.2 Double Deep Q-Network Architecture
@@
3.3 KAN
We now instantiate KAN as the DDQN head, following the background in Sec. 2-2.2.In our approach, the KAN-based architecture is structured with widths of [64, 8, 10] for the input, hidden, and output layers. Both the Q-network and target network use this compact configuration to capture complex feature interactions efficiently. Each KAN layer combines a spline function and a residual basis function, formulated as:
@@ 3.3 KAN
where $b(x)=silu(x)$ serves as a non-linear basis function, while $spline(x)=\sum_{i}c_{i}B_{i}(x)$ utilizes trainable B-splines for flexible approximation. With spline order $k=3$ over a specified grid, this layered structure dynamically adjusts to feature relevance throughout training. For optimization, $w_{b}$ and $w_{s}$ are initialized carefully: $w_{b}$ uses Xavier initialization, ensuring balanced layer activation scales, while $w_{s}$ is set to 1, with the initial spline function close to zero. The model’s spline grids adapt to changing input activations, extending KAN’s effective region for learning. The choice of KAN for the Q-network is theoretically motivated by its property of local plasticity, which mitigates catastrophic forgetting common in global MLP updates. We provide a detailed theoretical analysis of KAN approximation bias and gradient stability in Appendix 7.
3.4 MLP
The MLP-based architecture provides a more traditional setup for the DDQN framework, using a straightforward, fully connected design for both the Q-network and the target network. This configuration comprises two linear layers: a 64-neuron initial layer that maps the input features to a hidden space defined by a 32-neuron network width, followed by a ReLU activation, and a 10-neuron final linear layer that produces Q-values. The MLP requires a larger hidden layer with 32 neurons to approximate the complex mappings found in datasets, making it parameter-heavy relative to the KAN-based DDQN.
3.5 Training Procedure: KAN-based and MLP-based DDQN
The training procedure for the DDQN is illustrated in Algorithm 2 and Fig. 1, which outlines the main steps taken during each epoch. The algorithm initializes the environment and Q-networks and iteratively processes batches of experiences from a replay buffer $\mathcal{B}$. Each experience consists of the current state $s$, action $a$, reward $r$, and the next state $s^{\prime}$.
We employ two training algorithms to optimize our Q-network, each tailored to different model architectures: (1) a KAN-based training procedure that utilizes regularization for interpretability and stability, and (2) a standard MLP-based training procedure. Each training procedure integrates feedback cost when feature selection is enabled, promoting a minimalistic and interpretable feature set.
3.5.1 Temporal-Difference Target Computation
@@
$TD_{target}=r+\gamma\,Q_{\theta^{\prime}}\left(s^{\prime},\arg\max_{a}Q_{\theta}(s^{\prime},a)\right)$ (11)
@@
3.5.2 KAN-based Training with Regularization
@@
$\mathcal{L}_{\text{TD}}=L_{\delta}\left(Q_{\theta}(s,a)-TD_{target}\right)$ (12)
@@
Additionally, an entropy term is introduced to mitigate overconfidence in the predictions. The entropy $S(\Phi)$ for the KAN layer is given by:
@@
$\mathcal{L}_{\text{KAN}}=\mathcal{L}_{\text{TD}}+\lambda\mathcal{R}=L_{\delta}\left(Q_{\theta}(s,a),TD_{target}\right)+\lambda\left(\mathcal{R}_{\text{L1+Entropy}}+\mathcal{R}_{\text{Spline}}\right)$ (19)
3.5.3 Feedback Cost Integration for Feature Selection
@@
3.5.4 MLP-based Training
For the MLP-based training, we use the same TD target computation and $SmoothL1$ loss but omit the KAN-specific regularization terms. Like KAN-based training, feature selection introduces an additional feedback cost term to this loss: $\mathcal{L}_{\text{MLP, total}}=L_{\delta}\left(Q_{\theta}(s,a),TD_{target}\right)+\alpha\times\mathcal{L}_{\text{feedback}}$.
3.5.5 Optimization and Gradient Clipping
@@
3.6 Integration of HITL Feedback in DDQN
@@
3.6.1 Feature Selection Adjustment Using HITL Feedback
@@
3.6.2 Action-Selection Policy with HITL Feedback
@@
3.6.3 Iterative Learning Process
@@
3.6.4 Practical Implications and Robustness
While our framework utilizes simulated Gaussian feedback to model ideal expert attention, we discuss the practical implications, robustness to noisy annotations, and sensitivity to the loss function in Appendix 8.
4 Experimental Design
4.1 Dataset Preparation
We used a benchmark environment to standardize data preparation, feedback simulation, and visualizations across four datasets: MNIST, FashionMNIST, CIFAR-10, and CIFAR-100. Each dataset was resized to a standardized input dimension of $8\times 8$ pixels, with normalization applied based on empirically computed means and standard deviations specific to each dataset to aid in training convergence. For MNIST, we used a mean ($\mu$) of 0.1307 and a standard deviation ($\sigma_{SD}$) of 0.3081; for FashionMNIST, both $\mu$ and $\sigma_{SD}$ were set to 0.5. For CIFAR-10, normalization employed channel-wise $\mu$ of $(0.4914,0.4822,0.4465)$ and $\sigma_{SD}$ of $(0.247,0.243,0.261)$, while CIFAR-100 used $(0.5071,0.4865,0.4409)$ and $(0.2673,0.2564,0.2761)$ respectively. This normalization process adjusts pixel intensity $\mathbf{I}$ via the formula:
@@ 4.1 Dataset Preparation
4.2 Simulated Feedback Mechanism
@@ 4.3 Resources Used
4.4 Experimental Configurations
For the experiments conducted in this study, we employed two configurations for the MLP and KAN models, both set to a batch size of 128 and trained over 100 epochs. The configurations utilized a learning rate $1\times 10^{-3}$, a weight decay $1\times 10^{-4}$, and a discount factor ($\gamma$) of 0.99. The MLP configuration featured a width of 32 and an output size of 10 classes, while the KAN configuration used a width of 8 and a grid size of 3. Both models share a common input size of 64 features, a buffer size of 100,000, and a target update interval of 10. They incorporated warm-up episodes of 2 and utilized a beta distribution (see Table 1) for their stochastic feature selection processes, with an initial $\tau$ value set at 1.0. We strategically designed the architecture of our model by selecting the appropriate combination of convolutional layers and filters to ensure that the resulting selected feature map retains the same shape as the input and feedback size (64 for $8\times 8$ images). This choice is crucial, as it directly allows the simulated feedback (similar to input size) to correspond to the feature map dimensions, facilitating effective integration during the feature selection process. In this case, for 1, 2, and 3 convolutional layers, we must select 2, 4, and 8 filter sizes to maintain the desired selected feature shape. Detailed descriptions of the evaluation metrics (accuracy, calibration, sparsity, and complexity), training of baselines, ablation protocols, and visualization scope are provided in Appendices 9.1, 9.2, 9.3, and 9.4.
Figure 3: (a) Training accuracy per epoch comparison of KAN and MLP-based DDQN on MNIST (left) and FashionMNIST (right), and (b) confusion matrix of KAN-DDQN on MNIST (left) and FashionMNIST (right).
Figure 4: (a) Pruned KAN architecture after removing neurons with both low incoming and outgoing activation magnitudes, exposing a minimal functional skeleton; (b) feature importances extracted from auto-symbolic forms of learned splines, indicating how often input features appear in simplified expressions; (c) early training trajectory (first five steps), where the policy concentrates mass on discriminative inputs before spreading to additional features. Together, these link sparsity, symbolic structure, and policy behavior.
Figure 5: Activation functions for middle neurons in the KAN agent’s hidden layer on the MNIST dataset. Each subplot shows how a specific neuron transforms inputs via the spline activation, illustrating various activation behaviors across the layer.
Table 1: Performance of MLP-DDQN for different distributions on MNIST and FashionMNIST datasets. The $\mu$ and $\sigma_{SD}$ are computed over a 10-epoch moving window. Bold indicates the best performance.
@@ 4.4 Experimental Configurations
Table 2: Performance Metrics without Feature Selection on MNIST and FashionMNIST Datasets Using KAN and MLP Models.
@@
maP maR maF1 Test Accuracy
KAN MNIST 0.57 0.57 0.56 0.58 $\pm$ 0.01 5:10:50
MLP MNIST 0.45 0.51 0.47 0.52 $\pm$ 0.00 10:41
KAN FashionMNIST 0.63 0.65 0.63 0.64 $\pm$ 0.00 5:10:32
MLP FashionMNIST 0.55 0.60 0.57 0.59 $\pm$ 0.00 10:24
Table 3: Impact of FSNet parameters on test accuracy using MNIST and FashionMNIST datasets with feature selection
@@
Table 4: Impact of KAN Parameter Adjustments on MNIST and FashionMNIST Performance
@@ 5 Results
Our results reveal that KAN-DDQN outperforms MLP-DDQN in all cases on both MNIST and FashionMNIST datasets (see Fig. 3a and 3b) with 4 times fewer parameters and only $8\times 8$ sized images. Beyond top-1 accuracy, we report macro F1, macro AUC-PR, and calibration (ECE, NLL, Brier, with reliability diagrams) in Appendix LABEL:app:1.
5.1 Effect of Distribution Type in Feature Selection
@@
Table 1 demonstrates significant variability in MLP-DDQN performance across different distributions for both the MNIST and FashionMNIST datasets. We conducted this experiment to choose the best distribution that can mimic the human-like feature selection process. The beta distribution consistently yields the highest train and test accuracies on MNIST (79% and 78%, respectively), prompting us to conduct the remaining experiments with this distribution. Conversely, the Gumbel Softmax, Dirichlet, multinomial, Cauchy, and Laplace distributions perform poorly, with test accuracies as low as 12%. Notably, the uniform distribution also shows competitive results, particularly on the MNIST dataset, suggesting it can serve as a viable alternative for feature selection.
5.2 Effect of Feature Selection in Performance
5.2.1 Performance With Feature Selection
@@
Table 3 illustrates the performance of the KAN-DDQN and standard MLP-DDQN using feature selection across the MNIST and FashionMNIST datasets. In all cases, the hidden layer width was set to 32 neurons for the MLP and 8 for the KAN, reflecting the distinct architectures of each model. The KAN model consistently outperformed the MLP across all configurations, particularly on the MNIST dataset, achieving a test accuracy of 93% with a single convolutional layer, two filters, a width of 8, and a grid size of 3. For FashionMNIST, KAN similarly demonstrated superior test performance, reaching 83% accuracy under the same configuration, compared to MLP’s 74% accuracy. Across both datasets, KAN exhibited higher mean average precision (maP), recall (maR), and F1-score (maF1). For example, in the single-layer configuration on MNIST, KAN achieved scores of 0.92 for each metric, whereas MLP scored 0.84. The training time for the KAN model was generally longer due to its more intricate architecture, with the best-performing KAN model on MNIST requiring 4 hours and 35 minutes, compared to only 8 minutes for the MLP (see Table 3). Despite the increased computation, the accuracy gains suggest that KAN is efficient for applications where performance outweighs training time constraints. Notably, with the addition of a convolutional layer and increased filter size (e.g., 2 layers, 4 filters), KAN’s performance slightly declined, particularly on FashionMNIST, where test accuracy was 70%. This suggests a potential trade-off between model complexity and generalization, indicating that simpler KAN configurations, particularly those with a lower hidden layer width, may better retain discriminative power for smaller datasets.
5.2.2 Performance Without Feature Selection
@@
Table 2 provides a comparative view of model performance without feature selection, highlighting the impact of excluding feature selection on both models’ performance. Without feature selection, both KAN and MLP demonstrate considerably poor performance. Here, the KAN model achieves a notable decrease in test accuracy and maF1 compared with the MLP model. Specifically, on MNIST, KAN reaches a test accuracy of 58% with maF1 of 0.56, surpassing the MLP model’s 52% test accuracy and 0.47 maF1. On FashionMNIST, the KAN model continues to outperform, achieving a 64% test accuracy and 0.63 maF1, compared with MLP’s 59% accuracy and 0.57 maF1.
5.3 Effect of KAN Parameters in Performance
@@
Table 4 shows how KAN performance varies with parameter adjustments, focusing on width (number of neurons in the hidden layer) and grid size. Doubling the width of the feature maps from 8 to 16 yielded improvements in accuracy, especially on MNIST, where the test accuracy rose from 72% to 75%. However, this increase came at the expense of significantly longer training times, suggesting a trade-off between accuracy and computational efficiency. On FashionMNIST, similar parameter adjustments yielded more moderate gains. For example, increasing the grid size to 3 and the width to 16 raised test accuracy slightly from 71% to 72%. Moreover, altering the grid size from 2 to 3 produces varied outcomes: it preserves similar test accuracy in MNIST (0.72 to 0.73) but slightly decreases performance in FashionMNIST (0.71 to 0.70). This variation may stem from applying the B-spline in higher dimensions, warranting further exploration. This pattern suggests that KAN’s parameter sensitivity is dataset-dependent, with MNIST benefiting more from only width increases. Training time increased substantially with larger configurations. On MNIST, the KAN model with a grid size of 3 and a width of 16 took over 10 hours, while the smaller grid size of 2 and a width of 8 configurations was completed in 6 hours. Thus, reducing width and grid size may be advantageous for time-sensitive applications, especially where marginal accuracy improvements are not critical. So, larger hidden layer widths improve KAN’s accuracy but at the cost of increased training time. Balancing grid size and width for applications prioritizing efficiency may offer an optimal trade-off without sacrificing significant accuracy.
5.4 Cross-Dataset Performance and Scalability
A consolidated summary of results across MNIST, FashionMNIST, CIFAR-10, and CIFAR-100 is provided in Appendix 1010.1 and Table 5, highlighting consistent gains of KAN over MLP in accuracy, calibration, and scalability.
5.5 Ablation Studies
Ablation results are reported in Appendix 1010.2 and Table 6, confirming the critical role of instance-wise feature selection and differentiable gates in maintaining accuracy and calibration.
5.6 Complexity and Deployability Analysis
Model complexity, latency, and sparsity metrics are summarized in Appendix 1010.3 and Table 7, demonstrating that KAN remains deployable on edge devices despite moderate training overhead.
5.7 Interpretability of KAN-DDQN